OpenStudy (anonymous):

Find the first derivative for y=(x+1)^2 / 1+x^2

OpenStudy (anonymous):

quotient rule....

OpenStudy (anonymous):

with a chain rule

OpenStudy (anonymous):

I odnt get it =////

OpenStudy (anonymous):

dont*

OpenStudy (anonymous):

do you know the quotient rule?

OpenStudy (anonymous):

u= (x+1)^2 du/dx = 2(x+1) v= 1+x^2 dv/dx = 2x dy/dx = (inside-outside) / (bottom squared ) dy/dx = [ 2(x+1)(1+x^2) -2x(x+1)^2] / ( 1+x^2)^2 dy/dx = (x+1) [ 2(1+x^2) -2x(x+1) ] / (1+x^2)^2

OpenStudy (anonymous):

= (x+1)( 2 - 2x) / (1+x^2)^2

OpenStudy (anonymous):

\[= \frac{ 2(x+1)(1-x) }{(1+x^2)^2}\]

OpenStudy (anonymous):

\[= \frac{-2(x^2-1)}{(1+x^2)^2}\]

OpenStudy (anonymous):

Thanks mate